# LagCeresa.bib

### v1

@@ 7.3. Speculation for higher genus curves
$CH_{i}(J(C))\longrightarrow CH_{i}(J(C_{e})).$
If we choose a set of coordinates $(x_{1},\dots,x_{g})$ for $J(C)$ such that the tangent vector to $e$ in $J(C)$ is parallel to $(0,\dots,0,1)$, then this $g$-cycle is
@@
Let $Y$ be an oriented Lagrangian correspondence $Y\subset\overline{X}\times X^{\prime}$ between symplectic manifolds $X$ and $X^{\prime}$. Then $Y$ induces a well-defined map
$Y:$ $\hskip 2.84526ptalgCob(X)\longrightarrow algCob(X^{\prime})$
@@
Let $Y$ be an oriented Lagrangian correspondence $Y\subset\overline{X}\times X^{\prime}$ between symplectic manifold $X$ and $X^{\prime}$ which satisfies $\pi_{X}(Y)=X$. Then $Y$ induces a well-defined map
$Y:$ $\hskip 2.84526ptalgCob(X)\longrightarrow algCob(X^{\prime})$
@@
Let $L_{1}$ and $L_{2}$ be oriented Lagrangians in $X$. Assume there is an oriented Lagrangian correspondence
@@
$\varphi_{1324}:$ $\hskip 2.84526ptX\times X^{\prime}\times A\times A\longrightarrow X\times A\times X^{\prime}\times A$
@@
the map interchanging the factors. Then $\varphi_{1324}(Y\times\Delta_{A})$ is an oriented Lagrangian correspondence between $X\times A$ and $X^{\prime}\times A$. Define $\Gamma^{\prime}\subset X^{\prime}\times A$ as the image of $\Gamma$ under this correspondence. Observe
$\Gamma^{\prime}(F_{q})$ $=\pi_{X^{\prime}}(\Gamma^{\prime}\pitchfork(X^{\prime}\times F_{q}))=\pi_{X^{\prime}}(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(X^{\prime}\times F_{q}))$
$=\pi_{X^{\prime}}\left(\pi_{X^{\prime}\times A}(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(\Gamma\times X^{\prime}\times A)\pitchfork(X^{\prime}\times F_{q}))\right)$
$=\pi_{X^{\prime}}\left(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(\Gamma\times X^{\prime}\times F_{q})\right)=L^{\prime}_{2}.$
Similarly, we have $\Gamma^{\prime}(F_{p})=L_{1}^{\prime}$. Finally, we verify that $\pi_{B}(\Gamma^{\prime})=B$, in other words that $\forall b\in B$, $\Gamma^{\prime}(F_{b})$ is non-empty. This can be rephrased as the requirement that for all $b\in B$, there should exist an $x\in X$, an $x^{\prime}\in X^{\prime}$, and an $f\in F_{b}$ such that $(x,x^{\prime})\in Y$ and $(x,f)\in\Gamma$. We know that $\pi_{B}(\Gamma)=B$, therefore we are guaranteed the existence of some $x\in X$ and $f\in F_{b}$ such that $(x,f)\in\Gamma$. The existence of an associated $x^{\prime}\in X^{\prime}$ such that $(x,x^{\prime})\in Y$ follows from our condition that $\pi_{X}(Y)=X$. ∎
Here the condition that $\pi_{X}(Y)=X$ is used to ensure that $\pi_{B}(\Gamma^{\prime})=B$. Very generally, we expect that this condition can be relaxed while still yielding an equivalent definition for algebraic Lagrangian cobordisms. Recall from Section 3.1 that the classical notion of algebraic equivalence can be equivalently formulated in terms of cycles living in flat families over a curve, an abelian variety, or any smooth projective variety [parameter_spaces_alg_equiv]. It is reasonable to expect a similar statement tropically, and by extension for our definition of algebraic Lagrangian cobordisms to be equivalent to the existence of a Lagrangian correspondence $\Gamma^{\prime}$ supported on a tropical curve for instance, as in Lemma 3.7.
@@ 8.1. Tropical curves of hyperelliptic type
Recently, C. Ritter has shown [ceresa_period_tropical, Theorem A.] that, if $C$ is any genus $g$ tropical curve, the $\Omega^{g}_{1}$-flux obstruction for its Ceresa cycle $C-C^{-}$ vanishes if and only if $C$ is of hyperelliptic type. In particular, this implies that the symplectic flux vanishes for the Lagrangian Ceresa cycles associated to curves of hyperelliptic type. However, the mirror correspondence between rational equivalence and Lagrangian cobordisms suggests that they are not generally nullcobordant. These limitations of flux as an obstruction to Lagrangian cobordisms could already be noticed from the fact that it does not allow us to detect Lagrangians which are oriented cobordant, but not unobstructed cobordant.
@@ 8.2. Flux and characters of symplectic Torelli-type groups
A concept closely related to our notion of symplectic flux appeared in a paper by Reznikov [Reznikov_characteristic_classes_in_symplectic_topology, Section 4], where it is used to construct characters of a symplectic Torelli-type group $Symp_{0}(X)/Ham(X)$. Here $Symp_{0}(X)$ is the kernel of the map
$Symp(X)\longrightarrow\operatorname{Aut}(H_{odd}(X;\mathbb{Z})),$
and $Ham(X)\subset Symp(X)$ is the subgroup of Hamiltonian isotopies of $X$.
Given $f\in Symp_{0}(X)$ and $z\in H_{2k-1}(X;\mathbb{Z})$, define
@@
$\chi:Symp_{0}(X)/Ham(X)\longrightarrow\hom(H_{odd}(X;\mathbb{Z}),\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(X;\mathbb{Z})\right).$
Now let $X$ be a symplectic manifold of dimension $2(2k-1)$, and fix a class $z\in H_{2k-1}(X;\mathbb{Z})$ which admits a Lagrangian representative $L$. Then
$\chi(\cdot,z):Symp_{0}(X)/Ham(X)\longrightarrow\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(X;\mathbb{Z})\right)$
@@
${Symp_{0}(X)/Ham(X)}$${Cob^{or}(X)_{hom}}$${\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(X;\mathbb{Z})\right).}$$\scriptstyle{\chi(\cdot,z)}$$\scriptstyle{\varphi\mapsto L-\varphi(L)}$$\scriptstyle{\omega^{k}-flux}$
From this perspective, Theorem 7.1 can be translated into a non-triviality result for Reznikov’s character in $X(J(C))$. Indeed, recall that $L^{-}=\varphi^{-}(L)$ for some global symplectomorphism $\varphi^{-}$ which is in $Symp_{hom}(X)$. However, it is important to notice that the horizontal map from $Symp_{hom}(X)/Ham(X)$ to $Cob^{or}(X)_{hom}$ is not a group homomorphism, but a crossed homomorphism satisfying
$\varphi_{2}\circ\varphi_{2}\mapsto(L-\varphi_{1}(L))+(\varphi_{1}(L)-\varphi_{2}\circ\varphi_{1}(L)).$

### v2

@@ 6.4. Speculation for higher genus curves
$CH_{i}(J(C))\longrightarrow CH_{i}(J(C_{e}))$
between tropical Chow groups. If we choose a set of coordinates $(x_{1},\dots,x_{g})$ for $J(C)$ such that the tangent vector to $e$ in $J(C)$ is parallel to $(0,\dots,0,1)$, then this $g$-cycle is
@@
Let $Y$ be an oriented Lagrangian correspondence $Y\subset\overline{M}\times M^{\prime}$ between symplectic manifolds $M$ and $M^{\prime}$. Then $Y$ induces a well-defined map
$Y:$ $\hskip 2.84526ptalgCob(M)\longrightarrow algCob(M^{\prime})$
@@
Let $Y$ be an oriented Lagrangian correspondence $Y\subset\overline{M}\times M^{\prime}$ between symplectic manifold $M$ and $M^{\prime}$ which satisfies $\pi_{M}(Y)=M$. Then $Y$ induces a well-defined map
$Y:$ $\hskip 2.84526ptalgCob(M)\longrightarrow algCob(M^{\prime})$
@@
Let $L_{1}$ and $L_{2}$ be oriented Lagrangians in $M$. Assume there is an oriented Lagrangian correspondence
@@
$\varphi_{1324}:$ $\hskip 2.84526ptM\times M^{\prime}\times A\times A\longrightarrow M\times A\times M^{\prime}\times A$
@@
the map interchanging the factors. Then $\varphi_{1324}(Y\times\Delta_{A})$ is an oriented Lagrangian correspondence between $M\times A$ and $M^{\prime}\times A$. Define $\Gamma^{\prime}\subset M^{\prime}\times A$ as the image of $\Gamma$ under this correspondence. Observe
$\Gamma^{\prime}(F_{q})$ $=\pi_{M^{\prime}}(\Gamma^{\prime}\pitchfork(M^{\prime}\times F_{q}))=\pi_{M^{\prime}}(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(M^{\prime}\times F_{q}))$
$=\pi_{M^{\prime}}\left(\pi_{M^{\prime}\times A}(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(\Gamma\times M^{\prime}\times A)\pitchfork(M^{\prime}\times F_{q}))\right)$
$=\pi_{M^{\prime}}\left(\varphi_{1324}(Y\times\Delta_{A})\pitchfork(\Gamma\times M^{\prime}\times F_{q})\right)=L^{\prime}_{2}.$
Similarly, we have $\Gamma^{\prime}(F_{p})=L_{1}^{\prime}$. Finally, we verify that $\pi_{B}(\Gamma^{\prime})=B$, in other words that $\forall b\in B$, $\Gamma^{\prime}(F_{b})$ is non-empty. This can be rephrased as the requirement that for all $b\in B$, there should exist an $x\in M$, an $x^{\prime}\in M^{\prime}$, and an $f\in F_{b}$ such that $(x,x^{\prime})\in Y$ and $(x,f)\in\Gamma$. We know that $\pi_{B}(\Gamma)=B$, therefore we are guaranteed the existence of some $x\in M$ and $f\in F_{b}$ such that $(x,f)\in\Gamma$. The existence of an associated $x^{\prime}\in M^{\prime}$ such that $(x,x^{\prime})\in Y$ follows from our condition that $\pi_{M}(Y)=M$. ∎
Here the condition that $\pi_{M}(Y)=M$ is used to ensure that $\pi_{B}(\Gamma^{\prime})=B$. Very generally, we expect that this condition can be relaxed while still yielding an equivalent definition for algebraic Lagrangian cobordisms. Recall from Section 4.1 that the classical notion of algebraic equivalence can be equivalently formulated in terms of cycles living in flat families over a curve, an abelian variety, or any smooth projective variety [parameter_spaces_alg_equiv]. It is reasonable to expect a similar statement tropically, and by extension for our definition of algebraic Lagrangian cobordisms to be equivalent to the existence of a Lagrangian correspondence $\Gamma^{\prime}$ supported on a tropical curve for instance, as in Lemma 4.7.
@@ 7.1. Tropical curves of hyperelliptic type
Recently, C. Ritter has shown [ceresa_period_tropical, Theorem A.] that, if $C$ is any genus $g$ tropical curve, the $\Omega^{g}_{2,1}$-flux obstruction for its Ceresa cycle $C-C^{-}$ vanishes if and only if $C$ is of hyperelliptic type. In particular, this implies that the symplectic flux vanishes for the Lagrangian Ceresa cycles associated to curves of hyperelliptic type. On the other hand, given a nullhomologous Lagrangian $L$ in $M$ with “zero flux”, i.e. whose image $\Theta(L)$ in the Lefschetz Jacobian $LJ(M)$ introduced in Section 4.3 is zero, there is no easy way of
In fact, in cases where mirror symmetry holds, one would expect rational equivalence to be mirror to unobstructed (cylindrical) Lagrangian cobordism. We note that while flux detects when Lagrangians are not oriented nullcobordant, it may not detect when Lagrangians are oriented cobordant, but not unobstructed cobordant.
Classically in algebraic geometry, for a nullhomologous cycle to have “zero flux” is manifested in the fact that its image in the corresponding intermediate Jacobian is zero. However it is an open question whether However, the mirror correspondence between rational equivalence and Lagrangian cobordisms suggests that they are not generally nullcobordant. These limitations of flux as an obstruction to Lagrangian cobordisms could already be noticed from the fact that it does not allow us to detect Lagrangians which are oriented cobordant, but not unobstructed cobordant.
@@ 7.2. Flux and characters of symplectic Torelli-type groups
A concept closely related to our notion of symplectic flux appeared in a paper by Reznikov [Reznikov_characteristic_classes_in_symplectic_topology, Section 4], where it is used to construct characters of a symplectic Torelli-type group $Symp_{0}(M)/Ham(M)$. Here $Symp_{0}(M)$ is the kernel of the map
$Symp(M)\longrightarrow\operatorname{Aut}(H_{odd}(M;\mathbb{Z})),$
and $Ham(M)\subset Symp(M)$ is the subgroup of Hamiltonian isotopies of $M$.
Given $f\in Symp_{0}(M)$ and $z\in H_{2k-1}(M;\mathbb{Z})$, define
@@
$\chi:Symp_{0}(M)/Ham(M)\longrightarrow\hom(H_{odd}(M;\mathbb{Z}),\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(M;\mathbb{Z})\right).$
Now let $M$ be a symplectic manifold of dimension 2(2k1)2(2k-1), and fix a class $z\in H_{2k-1}(M;\mathbb{Z})$ which admits a Lagrangian representative $L$. Then
$\chi(\cdot,z):Symp_{0}(M)/Ham(M)\longrightarrow\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(M;\mathbb{Z})\right)$
@@
$Symp_{0}(M)/Ham(M)$$Cob^{or}(M)_{hom}$$\mathbb{R}/\left([\omega^{k}]\cdot H_{2k}(M;\mathbb{Z})\right).$$\scriptstyle{\chi(\cdot,z)}$$\scriptstyle{\varphi\mapsto L-\varphi(L)}$$\scriptstyle{\omega^{k}-flux}$
From this perspective, Theorem 6.1 can be translated into a non-triviality result for Reznikov’s character in $X(J(C))$. Indeed, recall that $L^{-}=\varphi^{-}(L)$ for some global symplectomorphism $\varphi^{-}$ which is in $Symp_{hom}(X(J(C)))$. However, it is important to notice that in general the horizontal map from $Symp_{hom}(M)/Ham(M)$ to $Cob^{or}(M)_{hom}$ is not a group homomorphism, but a crossed homomorphism satisfying
$\varphi_{2}\circ\varphi_{1}\mapsto(L-\varphi_{1}(L))+(\varphi_{1}(L)-\varphi_{2}\circ\varphi_{1}(L)).$